Chapter 09 Homework i Saved Required information [The following information applies to the questions displayed below.] Part 1Chapter 09 Homework i Saved 14 Required information The following information applies to the questions displayed below.] PartChapter 09 Homework i Saved Required information [The following information applies to the questions displayed below.) Part 3Chapter 09 Homework Saved Required information [The following information applies to the questions displayed below.) Part 4 oChapter 09 Homework i Saved Required information [The following information applies to the questions displayed below.] Part 55.2 Prepare journal entries for all the preceding transactions and events for 2018. (Do not round your intermediate calculati

Please help its a 5 part question and i dont know what to do
Answer #1
1) Maturity date
locust NBR fargo
date of the note 19-May 8-Jul 28-Nov
term of note 90 120 60
maturity date 17-Aug 5-Nov 27-Jan
2) interest due at maturity
principal * Rate * time = interest
locust 35,000 * 8% * 90/360 = 700
NBR 66,000 * 11% * 120/360 = 2420
Fargo 33,000 * 7% * 60/360 = 385
3) Amount in adjusting entry
Fargo Bank
principal * Rate * time = interest
interest to be acccrued in 2016 33,000 * 7% * 33/360 = 212
4) interest expense to be recorded in 2017
principal * Rate * time = interest
interest to recorded in 2018 33,000 * 7% * 27/360 = 173
Journal entries
Date Accounting titles & Explanations Debit Credit
2016
20-Apr inventory 37,500
Accounts payable 37,500
19-May Accounts payable 37,500
cash 2,500
notes payable 35,000
8-Jul Cash 66,000
notes payable 66,000
17-Aug notes payable 35,000
interest expense 700
cash 35,700
5-Nov notes payable 66,000
interest expense 2,420
cash 68,420
28-Nov Cash 33,000
notes payable 33,000
31-Dec interest expense 212
interest payable 212
2017
27-Jan notes payable 33,000
interest payable 212
interest expense 173
cash 33,385